### v1

# Optimal Trudinger-Moser inequalities on complete noncompact Riemannian manifolds: Revisit of the argument from the local inequalities to global ones

Abstract.
The main purpose of this short note, on the one hand, to is clarify some part of the proof of Theorem 1.3 in [8] in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones.
1. Introduction
The classical Pólya-Szegö symmetrization inequality plays a fundamental role in establishing geometric and functional inequalities in the Euclidean space. However, this symmetrization principle in general fails in many other non-Euclidean settings such as the Heisenberg group or Riemannian manifolds as well as on higher order Sobolev spaces even in the Euclidean space.
To circumvent this obstacle, Lam and Lu [5, 6] developed a symmetrization-free method to prove critical global Trudinger-Moser inequalities on the entire Heisenberg group, and Adams inequalities on the higher order Sobolev spaces on the entire Euclidean spaces. Lam, Lu and Tang [7] establish the subcritical Trudinger-Moser inequality on the Heisenberg group. The essence of these techniques is to establish the local inequalities on finite domains and then pass them to global inequalities using an argument of level set of functions under consideration. This kind of argument avoids using the Póly-Szegö inequality and thus is also effective in proving the concentration-compactness principle in settings where such symmtrization principle is absent ([10, 11]). This argument has been extensively used by a number of authors and we only refer the reader to see also e.g. [14], [2], [12, 13] for adapations of this circle of ideas.
This sort of ideas has also been used in our proofs of the critical Trudinger-Moser inequality (Theorem 1.3) on complete and noncompact Riemannian manifolds [8]. In the proof of Theorem 1.3 in [8], some clarifications need to be made about the uniformity of the upper bound of the Trudinger-Moser inequality in terms of the bounded domains of the level sets involved. To avoid the complexity, we will then simply give another simple proof by only establishing the uniformity of the bounds of the Trudinger-Moser inequality on geodesic balls within the manifolds (still going through the same proof of Theorem 3.1 in [8] on the geodesic balls or simply applying Theorem 3.1 in [8] on geodesic balls where the Green functions have uniform estimates) together with a Gromov’s covering lemma (similar to the proof of subcritical case in [8]). The proof of Theorem 3.1 in [8] remains valid for geodesic balls with uniform constants for the Trudinger-Moser inequalities for such balls independent of the location of the balls. Since the main purpose of this short note is to clarify the proof of Theorem 1.3 in [8] and in the meantime to give an alternative proof from the level set argument passing from the local inequalities to global one, we have chosen not to give a comprehensive account of the subject (see e.g. [8] for relevant references in the literature).
2. Some useful known lemmas
We recall the critical Moser-Trudinger inequality in [8].
Theorem 2.1.
Let $(M,g)$ be a complete noncompact Riemannian manifold whose Ricci curvature is lower bounded, i.e. $\textit{Ric}\geq\lambda g$ for some constant $\lambda\in\mathbb{R}$. Moreover, we assume that its injectivity radius has a lower bound, i.e. $\textit{inj}(M,g)\geq i>0$ for some constant $i>0$. Then for every $\tau>0$, there exists a constant $C=C(n,\tau,M)$ such that
[…]
Moreover, $\alpha_{n}$ is sharp.
The Ricci lower bound condition guarantees a finite overlapping covering of $M$. This is the so-called Gromov’s covering lemma, which states as follows:
Lemma 2.1 (Hebey [4], Lemma 1.6).
Let $(M,g)$ be a complete Riemannian manifold with Ricci lower bound $\lambda$, let $\rho>0$ be given, then there exists a sequence of points $\{x_{j}\}$ on $M$ such that for any $r\geq\rho$:
  1. (1)
    the family $\{B_{r}(x_{j})\}$ is a uniformly locally finite covering of $M$, there exists an upper bound for the overlapping number in terms of $n,\rho,r,\lambda$;
  2. (2)
    for any $i\neq j$, $B_{r/2}(x_{i})\cap B_{r/2}(x_{j})=\emptyset$.
On the other hand, our geometric assumptions guarantee the existence of harmonic coordinates whose harmonic radius is uniformly bounded from below. To be precise, we have
Lemma 2.2 (Hebey [4], Theorem 1.3).
Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy our assumptions. $\Omega(\delta)=\{x\in M:\textit{dist}(x,\Omega)<\delta\}$. Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in\Omega$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
Combining Lemma 2.1 and Lemma 2.2, we obtain a sequence of balls $\{B_{\delta}(x_{j})\}$ satisfies following:
  • For $Q>1$ fixed (close to 1), $\delta\leq r_{H}(Q,0,\alpha)$ and $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ and moreover, $g_{ij}$ satisfies the $C^{0,\alpha}$ estimate:
    $r_{H}^{\alpha}\sup_{y\neq z}\frac{|g_{ij}(z)-g_{ij}(y)|}{d(y,z)^{\alpha}}\leq Q-1.$
  • $\{B_{\delta/2}(x_{j})\}$ covers $M$ and when $i\neq j$, $B_{\delta/4}(x_{i})\cap B_{\delta/4}(x_{j})=\emptyset$;
  • There exists $N=N(n,\lambda,\delta)$ such that any $x\in M$ belongs to at most $N$ of $B_{\delta}(x_{j})$’s.
3. Proof of Theorem 2.1
To give another proof of Theorem 2.1, we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one. We first prove
Theorem 3.1.
Let $B_{\delta}(x_{j})$ be any ball mentioned above, then there exists a constant $C=C(n,\lambda,i)$ such that
[…]
The proof of Theorem 3.1 consists of 2 steps:
Step 1: Since $u\in W^{1,n}_{0}(B_{\delta}(x_{j}))$, we recall the pointwise representation formula of $u(x)$ in terms of the Dirichlet Green’s function:
$|u(x)|\leq\int_{B_{\delta}(x_{j})}|\nabla G(x,y)||\nabla u(y)|dV_{g}(y).$
The existence of the Green’s function comes from a classical paramatrix argument (see Aubin, Nonlinear Analysis on Manifolds. Monge-Ampére Equations, 1982, Chapter 4). To be precise, for some positive integer $k>\frac{n}{2}$,
$G(x,y)=H(x,y)+\sum_{l=1}^{k}\int_{B_{\delta}(x_{j})}\Gamma_{l}(x,z)H(z,y)dV_{g}(z)+F(x,y),$
where $H(x,y)=\frac{1}{(n-2)\omega_{n-1}}d(x,y)^{2-n}f(d(x,y))$ is the fundamental solution of Laplacian multiplied by a smooth cutoff function $f$, which is identically equal to 1 near diagonal and vanishes when $d(x,y)\geq\textit{inj}(M,g)$. $\Gamma(x,y)=\Gamma_{1}(x,y)=-\Delta_{y}H(x,y)$ and we inductively define
$\Gamma_{l+1}(x,y)=\int_{B_{\delta}(x_{j})}\Gamma_{l}(x,z)\Gamma(z,y)dV_{g}(z).$
The error term $F$ satisfies the equation
(3.1) $\begin{cases}\Delta_{y}F(x,y)=\Gamma_{k+1}(x,y),\\ \ F(x,y)=0\ \ \ \textit{for }y\in\partial B_{\delta}(x_{j}).\end{cases}$
We claim that
(3.2) $|\nabla_{y}G(x,y)|\leq\frac{1}{\omega_{n-1}}d(x,y)^{1-n}(1+Cd(x,y)),$
where $C=C(n,\lambda,Q,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
$\int_{B_{\delta}(x_{j})}\Gamma(x,z)H(z,y)dV_{g}(z).$
From the asymptotic behavior near singularity, we need to prove the following lemma:
Lemma 3.1.
For any $\alpha,\beta>0$ such that $\alpha+\beta<n$, there exists a constant $C=C(n,\alpha,\beta,Q)$ such that
$\int_{B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)\leq Cd(x,y)^{\alpha+\beta-n}.$
Proof.
Let $\rho=\frac{d(x,y)}{2}$. Then
(3.3) $\int_{B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
(3.4) $=$ $\left(\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}+\int_{(B_{3\rho}(y)\setminus B_{\rho}(x))\cap B_{\delta}(x_{j})}+\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}\right)d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
For the first term of the right hand side,
$\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
$\leq$ $\rho^{\beta-n}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}dV_{g}(z),$
$\leq$ $Q^{n/2}\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
$\leq$ $Q^{n/2}\omega_{n-1}\rho^{\alpha+\beta-n},$
where in the third line we use the property of harmonic coordinates $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$. Similarly we can estimate the second term of the right hand side of (3.3). For the third term, we have
(3.5) $\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
(3.6) $\leq$ $C_{n}\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(z,y)^{\alpha+\beta-2n}dV_{g}(z),$
(3.7) $\leq$ $C_{n}Q^{n/2}\omega_{n-1}\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$
(3.8) $\leq$ $C_{n,\alpha,\beta}Q^{n/2}\omega_{n-1}((3\rho)^{\alpha+\beta-n}-\delta^{\alpha+\beta-n}),$
(3.9) $\leq$ $C_{n,\alpha,\beta}Q^{n/2}\omega_{n-1}\rho^{\alpha+\beta-n}.$
From Lemma 3.1, $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a constant $C=C(n,\lambda,i,Q)$. Therefore, to obtain the uniform estimate for Green’s function, we need to prove the uniform $C^{1}$ estimate of $F(x,y)$.
Lemma 3.2.
Let $F(x,y)$ be a $W^{1,2}_{0}(B_{\delta}(x_{j}))$ solution of the Dirichlet problem (3.1), there exists a positive constant $C=C(n,\lambda,i,Q)$ such that $\nabla_{y}F(x,y)||_{\infty}\leq C$.
Proof.
Since $\Gamma_{k+1}$ in (3.1) has no singularity, it is smooth and hence $F$ is the classical smooth solution. We first establish the $C^{0}$ estimate of $F(x,y)$, this follows from a classical barrier argument: consider the following Dirichlet problem on $B_{\delta}(x_{j})$:
(3.10) $\begin{cases}-\Delta u=f(x),\ \ \ &x\in B_{\delta}(x_{j});\\ u=0,\ \ \ &x\in\partial B_{\delta}(x_{j}).\end{cases}$
We are to show: there exists a constant $C=C(n,Q,\delta)$ such that
[…]
In fact, let w(x)=u(x)z(x)w(x)=u(x)-z(x), where
[…]
$x$ is the position variable in terms of harmonic coordinates, and the constant $A$ is left to be chosen. It is easy to verify that
[…]
Since with harmonic coordinates, we have
$\Delta|x|^{2}=2|\nabla x|^{2}+x\Delta x\in[2nQ^{-1},2nQ],$
then we can set $A=2nQ^{-1}$ and hence $\Delta w\leq 0$. Since $w\leq 0$ on $\partial B_{\delta}(x_{j})$, by maximum principle, we have
[…]
Repeat the above argument for u(x)-u(x) then we have
[…]
Therefore we obtain the $C^{0}$ estimate of $F$. Now we recall the Schauder interior estimate of elliptic equations (see Gilbarg-Trudinger [3], Theorem 6.2):
Theorem 3.2 (Schauder estimates).
Let $u\in C^{2,\alpha}(\Omega)$ be a bounded solution in $\Omega$ and the equation
$a^{ij}\partial_{ij}u+b^{i}\partial_{i}u+cu=f,$
where $f\in C^{\alpha}$ and there are positive constants $C_{1},C_{2}$ such that $a^{ij}\xi_{i}\xi_{j}\geq C_{1}|\xi|^{2}$ and
$|a^{ij}|_{C^{0,\alpha}},\ |b^{i}|^{(1)}_{C^{0,\alpha}},\ |c|^{(2)}_{C^{0,\alpha}}\leq C_{2}.$
[…]
In our case, $b^{i},c$ vanish and from Theorem 2.2 in harmonic coordinates $a^{ij}=g^{ij}$ belongs to $C^{0,\alpha}$. $f$ corresponds to $\Gamma_{k}$, which is proved to be uniformly bounded from Lemma 3.1. Therefore we have $|F|^{*}_{C^{2,\alpha}}\leq C(n,\lambda,i,Q,\alpha)$. In specific, $\nabla_{y}F(x,y)||_{\infty}\leq C(n,\lambda,i,Q)$ and we complete the proof. ∎
Step 2: We use Adams’ method to obtain the equivalent version of Theorem 3.1:
Theorem 3.3.
Let $B_{\delta}(x_{j})$ be defined as previous and $T$ be the convolution operator such that
$Tu(x)=\int_{B_{\delta}(x_{j})}|\nabla G(x,y)||\nabla u(y)|dV_{g}(y).$
Then there exists a constant $C=C(n,\lambda,i,Q,\textit{Vol}(B_{\delta}(x_{j})))$, which is increasing with respect to $\textit{Vol}(B_{\delta}(x_{j}))$, such that
[…]
The proof of Theorem 3.3 is exactly the same as that of Theorem 3.1 in [8] which is one of the main results. The parameters $\lambda,i$ do not explicitly appear in the estimates, but will determine $Q$ and $r_{H}(\alpha,0,Q)$, which gives the bounds of $\delta$. This shows the constants in the Moser-Trudinger inequality on any geodesic balls only depend on the radius $\delta$ which in turn depends on $\lambda$ and $i$.
Proof of Theorem 2.1.
For each $j$, we take cutoff function $\psi_{j}\in C_{0}^{\infty}(B_{\delta}(x_{j}))$ satisfying $0\leq\psi_{j}\leq 1$, $\psi_{j}=1$ on $B_{\frac{\delta}{2}}(x_{j})$ and $|\nabla\psi_{j}|\leq\frac{4}{\delta}$. Then $|\nabla\psi_{j}^{2}|=2\psi_{j}|\nabla\psi_{j}|\leq\frac{8}{\delta}\psi_{j}$ and $1\leq\sum_{j}\psi_{j}(x)\leq N$. Thus we have
(3.11) $\int_{M}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g}$
(3.12) $\leq$ $\sum_{j}\int_{B_{\delta/2}(x_{j})}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g},$
(3.13) $\leq$ $\sum_{j}\int_{B_{\delta}(x_{j})}\phi_{n}(\alpha_{n}|\psi_{j}^{2}u|^{\frac{n}{n-1}})dV_{g},$
[…]
(3.16) $\leq$ $C(n,\lambda,i)2^{n}\sum_{j}\int_{M}\psi_{j}^{2n}|\nabla u|^{n}+|\nabla\psi_{j}^{2}|^{n}|u|^{n}dV_{g},$
(3.17) $\leq$ $C(n,\lambda,i)2^{n}\left(N\int_{M}|\nabla u|^{n}dV_{g}+\frac{8^{n}}{\delta^{n}}N\int_{M}|u|^{n}dV_{g}\right),$
(3.18) $\leq$ $C(n,\lambda,i)(N+N\frac{8^{n}}{\delta^{n}\tau^{n}}).$
This is a constant only depending on the initial geometry of the manifold.

### v2

@@
 Abstract.
-The main purpose of this short note, on the one hand, to is clarify some part of the proof of Theorem 1.3 in [8] in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones.
+The main purpose of this short note, on the one hand, is to clarify some part of the proof of Theorem 1.3 in [10] using the level set argument in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones by using Gromov’s covering lemma.
+Moreover, we show that the Green functions for the Dirichlet problem on any small geodesic balls (or harmonic coordinate charts) on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls (or the charts).
+This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set and thus rigorize the proof of Theorem 1.3 in [10].
+This latter result of Green’s function estimates is of its independent interest and can be found useful elsewhere.
 1. Introduction
@@
 [16], [4], [14, 15] for adapations of this circle of ideas.
-This sort of ideas has also been used in our proofs of the critical Trudinger-Moser inequality (Theorem 1.3) on complete and noncompact Riemannian manifolds [8].
-In the proof of Theorem 1.3 in [8], some clarifications need to be made about the uniformity of the upper bound of the Trudinger-Moser inequality in terms of the bounded domains of the level sets involved.
-To avoid the complexity, we will then simply give another simple proof by only establishing the uniformity of the bounds of the Trudinger-Moser inequality on geodesic balls within the manifolds (still going through the same proof of Theorem 3.1 in [8] on the geodesic balls or simply applying Theorem 3.1 in [8] on geodesic balls where the Green functions have uniform estimates) together with a Gromov’s covering lemma (similar to the proof of subcritical case in [8]).
-The proof of Theorem 3.1 in [8] remains valid for geodesic balls with uniform constants for the Trudinger-Moser inequalities for such balls independent of the location of the balls.
-Since the main purpose of this short note is to clarify the proof of Theorem 1.3 in [8] and in the meantime to give an alternative proof from the level set argument passing from the local inequalities to global one, we have chosen not to give a comprehensive account of the subject (see e.g.
-[8] for relevant references in the literature).
-2. Some useful known lemmas
-We recall the critical Moser-Trudinger inequality in [8].
+This sort of ideas has also been used in our proofs of the critical Trudinger-Moser inequality (Theorem 1.3) on complete and noncompact Riemannian manifolds [10]. To be precise, we established
 Theorem 1.1.
@@
 Moreover, $\alpha_{n}$ is sharp.
-The Ricci lower bound condition guarantees a finite overlapping covering of $M$. This is the so-called Gromov’s covering lemma, which states as follows:
-Lemma 2.1 (Hebey [4], Lemma 1.6).
-Let $(M,g)$ be a complete Riemannian manifold with Ricci lower bound $\lambda$, let $\rho>0$ be given, then there exists a sequence of points $\{x_{j}\}$ on $M$ such that for any $r\geq\rho$:
+In the proof of Theorem 1.1 some clarifications need to be made about the uniformity of the upper bound of the Trudinger-Moser inequality in terms of the bounded domains of the level sets involved.
+In this note, we will then simply give another simple proof by only establishing the uniformity of the bounds of the Trudinger-Moser inequality on geodesic balls within the manifolds (still going through the same proof of Theorem 3.1 in [10] on the geodesic balls or simply applying Theorem 3.1 in [10] on geodesic balls where the Green functions have uniform estimates) together with a Gromov’s covering lemma (similar to the proof of subcritical case in [10]).
+The proof of Theorem 3.1 in [10] remains valid for geodesic balls with uniform constants for the Trudinger-Moser inequalities for such balls independent of the location of the balls.
+The main result of this note is that the Green functions for the Dirichlet problem on any small geodesic balls on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls.
+This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set.
+More precisely, in the proof of Theorem 1.3 in [10], the level set $\Omega(u)$ can be covered by a finite number of small geodesic balls (or harmonic coordinate charts) and the number of such balls (or charts) is uniformly bounded.
+Therefore, the same level argument from local inequalities to the global ones in [10] are justified.
+Since the main purpose of this short note is to rigorize the proof of Theorem 1.3 in [10] and in the meantime to give an alternative proof from the level set argument passing from the local inequalities to global one, we have chosen not to give a comprehensive account of the subject (see e.g.
+[10] for relevant references in the literature).
+2. Some preliminaries in differential geometry
+We first recall the Gromov’s covering lemma:
+Lemma 2.1 (Lemma 1.6 of Hebey [6]).
+Let $(M,g)$ be a complete Riemannian manifold with Ricci lower bound $\lambda$, let $\delta>0$ be given, then there exists a sequence of points $\{x_{j}\}$ on $M$ such that for any $r\geq\delta$:
   1. (1)
-    the family $\{B_{r}(x_{j})\}$ is a uniformly locally finite covering of $M$, there exists an upper bound for the overlapping number in terms of $n,\rho,r,\lambda$;
+    the family $\{B_{r}(x_{j})\}$ is a uniformly locally finite covering of $M$, there exists an upper bound for the overlapping number in terms of $n,\delta,r,\lambda$;
   2. (2)
     for any $i\neq j$, $B_{r/2}(x_{i})\cap B_{r/2}(x_{j})=\emptyset$.
-On the other hand, our geometric assumptions guarantee the existence of harmonic coordinates whose harmonic radius is uniformly bounded from below. To be precise, we have
-Lemma 2.2 (Hebey [4], Theorem 1.3).
-Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy our assumptions.
-$\Omega(\delta)=\{x\in M:\textit{dist}(x,\Omega)<\delta\}$.
+On the other hand, lower bounds for both Ricci curvature and injectivity radius guarantee the existence of $C^{0,\alpha}$ harmonic coordinates.
+This is due to Anderson-Cheeger [1] (see also Hebey [6]) and to be precise:
+Lemma 2.2.
+Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy $\textit{Ric}\geq\lambda g$ and $\textit{inj}(M,g)\geq i>0$.
 Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in\Omega$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
@@
     $r_{H}^{\alpha}\sup_{y\neq z}\frac{|g_{ij}(z)-g_{ij}(y)|}{d(y,z)^{\alpha}}\leq Q-1.$
+  • In the overlapping charts $B_{\delta}(x_{i})\cap B_{\delta}(x_{j})$, the transition function satifies
+    $y^{k}\circ(x^{l})^{-1}||_{C^{1,\alpha}}\leq C(n,Q,\lambda,\alpha,\delta,i).$
   • $\{B_{\delta/2}(x_{j})\}$ covers $M$ and when $i\neq j$, $B_{\delta/4}(x_{i})\cap B_{\delta/4}(x_{j})=\emptyset$;
   • There exists $N=N(n,\lambda,\delta)$ such that any $x\in M$ belongs to at most $N$ of $B_{\delta}(x_{j})$’s.
-3. Proof of Theorem 2.1
-To give another proof of Theorem 2.1, we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one.
-We first prove
-Theorem 3.1.
-Let $B_{\delta}(x_{j})$ be any ball mentioned above, then there exists a constant $C=C(n,\lambda,i)$ such that
+Remark 2.1.
+Lemma 2.2 is due to Anderson-Cheeger [1].
+The $C^{0,\alpha}$ result was achieved by the existence of $W^{1,p}$- harmonic coordinates and a classical Sobolev embedding for $\alpha=1-\frac{n}{p}$.
+This means in each local chart, the metric tensor satisfies
 […]
-The proof of Theorem 3.1 consists of 2 steps:
-Step 1: Since $u\in W^{1,n}_{0}(B_{\delta}(x_{j}))$, we recall the pointwise representation formula of $u(x)$ in terms of the Dirichlet Green’s function:
+Similar embedding also appears for the coordinate functions $x^{i},i=1,\dots,n$ and we have $W^{2,p}$ uniform bounds.
+3. Green’s function estimate
+For $u\in W^{1,n}_{0}(B_{\delta}(x_{j}))$, we recall the pointwise representation formula of $u(x)$ in terms of the Dirichlet Green’s function:
 $|u(x)|\leq\int_{B_{\delta}(x_{j})}|\nabla G(x,y)||\nabla u(y)|dV_{g}(y).$
-The existence of the Green’s function comes from a classical paramatrix argument (see Aubin, Nonlinear Analysis on Manifolds.
-Monge-Ampére Equations, 1982, Chapter 4).
-To be precise, for some positive integer $k>\frac{n}{2}$,
+The existence of the Green’s function comes from a classical paramatrix argument (see e.g. Chapter 4 of Aubin [2]). To be precise, for some positive integer $k>\frac{n}{2}$,
 $G(x,y)=H(x,y)+\sum_{l=1}^{k}\int_{B_{\delta}(x_{j})}\Gamma_{l}(x,z)H(z,y)dV_{g}(z)+F(x,y),$
@@
 (3.2) $|\nabla_{y}G(x,y)|\leq\frac{1}{\omega_{n-1}}d(x,y)^{1-n}(1+Cd(x,y)),$
-where $C=C(n,\lambda,Q,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
+where $C=C(n,\lambda,Q,\lambda,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
 $\int_{B_{\delta}(x_{j})}\Gamma(x,z)H(z,y)dV_{g}(z).$
@@
 (3.4) $=$ $\left(\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}+\int_{(B_{3\rho}(y)\setminus B_{\rho}(x))\cap B_{\delta}(x_{j})}+\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}\right)d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
-For the first term of the right hand side,
+It is well known that in polar coordinates
+$dV=r^{n-1}\left(1-\frac{\textit{Ric}(\theta)}{3}r^{2}+o(r^{2})\right)d\theta dr.$
+When $\delta$ is small, providing Ricci curvature lower bounded, the first term of the right hand side,
 $\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
 $\leq$ $\rho^{\beta-n}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}dV_{g}(z),$
-$\leq$ $Q^{n/2}\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
-$\leq$ $Q^{n/2}\omega_{n-1}\rho^{\alpha+\beta-n},$
-where in the third line we use the property of harmonic coordinates $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$.
-Similarly we can estimate the second term of the right hand side of (3.3).
-For the third term, we have
+$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
+$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\alpha+\beta-n},$
+Similarly we can estimate the second term of the right hand side of (3.3). For the third term, we have
 (3.5) $\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
 (3.6) $\leq$ $C_{n}\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(z,y)^{\alpha+\beta-2n}dV_{g}(z),$
-(3.7) $\leq$ $C_{n}Q^{n/2}\omega_{n-1}\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$
-(3.8) $\leq$ $C_{n,\alpha,\beta}Q^{n/2}\omega_{n-1}((3\rho)^{\alpha+\beta-n}-\delta^{\alpha+\beta-n}),$
-(3.9) $\leq$ $C_{n,\alpha,\beta}Q^{n/2}\omega_{n-1}\rho^{\alpha+\beta-n}.$
-From Lemma 3.1, $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a constant $C=C(n,\lambda,i,Q)$.
-Therefore, to obtain the uniform estimate for Green’s function, we need to prove the uniform $C^{1}$ estimate of $F(x,y)$.
+(3.7) $\leq$ $C_{n}C(\lambda,\delta)\omega_{n-1}\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$
+(3.8) $\leq$ $C_{n,\alpha,\beta,\lambda,\delta}\omega_{n-1}((3\rho)^{\alpha+\beta-n}-\delta^{\alpha+\beta-n}),$
+(3.9) $\leq$ $C_{n,\alpha,\beta,\lambda,\delta}\omega_{n-1}\rho^{\alpha+\beta-n}.$
+It is easy to verify that in polar coordinates,
+[…]
+where $r=d(x,y)$ and $|\widetilde{g}|=\textit{det}(\widetilde{g}_{ij})$ denotes the determinant of the corresponding metric tensor.
+By carefully choosing the smooth cut-off function $f$, both the first two terms $|(n-3)f^{\prime}-rf^{\prime\prime}|$ and the coefficient of the third term $((n-2)f-rf^{\prime})$ can be uniformly bounded by $C(n,\delta,i)$.
+With the help of Lemma 3.1, we only need to consider the integral of the form
+$\int_{B_{\delta}(x_{j})}|\nabla_{z}\ln\sqrt{|\widetilde{g}|}|(z)H(z,y)dV_{z}.$
+An easy change of variablse from polar coordinates to harmonic coordinates shows
+$\widetilde{g}_{ij}=g_{kl}\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\frac{\partial x^{l}}{\partial\widetilde{x}^{j}},$
+where $\{x^{k}\}$ and $\{\widetilde{x}^{i}\}$ denotes the coordinate variables respectively.
+For simplicity we denote the transition matrix $A=\left(\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\right)$ then
+$|\nabla\ln\sqrt{|\widetilde{g}|}|\leq\frac{|\nabla|g||}{2|g|}+\frac{|\nabla|A||}{|A|}.$
+Since in harmonic coordinates $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$, we have
+$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|g||}{|g|}(z)H(z,y)dV_{z}\right|$ $\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$
+Similar to the argument in the proof of Lemma 3.1, the integral can be splitted into 3 pieces.
+Here we just check the part over $B_{\rho}(x)\cap B_{\delta}(x_{j})$, the other two parts can be treated in a similar manner:
+$\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$ $\leq\frac{C(n)}{d(x,y)^{n-2}}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}|\nabla g|dV_{z};$
+[…]
+$\leq\frac{C(n,Q)}{d(x,y)^{n-2-\frac{n}{p^{\prime}}}}.$
+On the other hand, $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ implies there exists $C(n,Q)>0$ such that $|A|\geq C(n,Q)>0$. Hence we have
+$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|A||}{|A|}(z)H(z,y)dV_{z}\right|\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla|A|(z)|}{d(z,y)^{n-2}}dV_{z}$
+We can estimate the above integral in a similar manner, but this time we make use of the $W^{2,p}$ estimate of the harmonic coordinate function $x^{i},i=1,\dots,n$.
+Indeed, in the derivation of the existence of harmonic coordinates, the $L^{p}$ theory of elliptic operators gives the uniform bound of $x^{i}||_{W^{2,p}}$, where $p$ satisfies $\alpha=1-\frac{n}{p}$.
+Hence there exists $C(n,\lambda,i,Q)$ such that $||\nabla A|_{L^{p}}\leq C(n,\lambda,i,Q)$.
+Therefore from Lemma 3.1, $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a constant $C=C(n,\lambda,i,Q)$.
+To obtain the uniform estimate for Green’s function, it suffices to prove the uniform $C^{1}$ estimate of $F(x,y)$.
 Lemma 3.2.
@@
 […]
-Therefore we obtain the $C^{0}$ estimate of $F$. Now we recall the Schauder interior estimate of elliptic equations (see Gilbarg-Trudinger [3], Theorem 6.2):
-Theorem 3.2 (Schauder estimates).
-Let $u\in C^{2,\alpha}(\Omega)$ be a bounded solution in $\Omega$ and the equation
-$a^{ij}\partial_{ij}u+b^{i}\partial_{i}u+cu=f,$
-where $f\in C^{\alpha}$ and there are positive constants $C_{1},C_{2}$ such that $a^{ij}\xi_{i}\xi_{j}\geq C_{1}|\xi|^{2}$ and
-$|a^{ij}|_{C^{0,\alpha}},\ |b^{i}|^{(1)}_{C^{0,\alpha}},\ |c|^{(2)}_{C^{0,\alpha}}\leq C_{2}.$
+Therefore we obtain the $C^{0}$ estimate of $F$.
+Now we proceed the $C^{1,\alpha}$ Schauder estimate on $B_{\delta}(x_{j})$ so we obtain the $L^{\infty}$ estimate of $\nabla_{y}F(x,y)$.
+The proof reduces the problem to Euclidean boundary $C^{1,\alpha}$ estimates via harmonic coordinates and a flattening of the boundary.
+The argument is classical and standard, we only sketch the proof.
+Step 1: Reduction to a Euclidean domain.
+Let $\Phi:B_{r}(p)\to\Omega\subset\mathbb{R}^{n}$ be the harmonic coordinate chart.
+Let $\widetilde{u}=u\circ\Phi^{-1}$, $\widetilde{f}=f\circ\Phi^{-1}$, and $\widetilde{\varphi}=\varphi\circ\Phi^{-1}$.
+Then $\widetilde{u}$ solves a uniformly elliptic equation in divergence form on $\Omega$:
 […]
-In our case, $b^{i},c$ vanish and from Theorem 2.2 in harmonic coordinates $a^{ij}=g^{ij}$ belongs to $C^{0,\alpha}$.
-$f$ corresponds to $\Gamma_{k}$, which is proved to be uniformly bounded from Lemma 3.1.
-Therefore we have $|F|^{*}_{C^{2,\alpha}}\leq C(n,\lambda,i,Q,\alpha)$.
-In specific, $\nabla_{y}F(x,y)||_{\infty}\leq C(n,\lambda,i,Q)$ and we complete the proof.
-∎
-Step 2: We use Adams’ method to obtain the equivalent version of Theorem 3.1:
+with boundary data $\widetilde{u}=\widetilde{\varphi}$ on $\partial\Omega$. The coefficients satisfy
+$\widetilde{a}^{ij},\widetilde{b}^{i},\widetilde{c}\in C^{\alpha}(\overline{\Omega}),\qquad\|\widetilde{a}^{ij}\|_{C^{\alpha}}+\|\widetilde{b}^{i}\|_{C^{\alpha}}+\|\widetilde{c}\|_{C^{\alpha}}\leq C(n,K,i_{0},r).$
+Step 2: Flattening the boundary.
+Since $r<i_{0}$, the boundary $\partial B_{r}(p)$ is smooth, and thus $\partial\Omega$ is $C^{1,\alpha}$.
+Cover $\partial\Omega$ by finitely many coordinate patches $\{U_{k}\}$ such that each patch is mapped by a $C^{1,\alpha}$-diffeomorphism $\Psi_{k}$ to a half-ball $B_{1}^{+}\subset\mathbb{R}^{n}$.
+Let $\{\eta_{k}\}$ be a partition of unity subordinate to this cover.
+Step 3: Apply Euclidean boundary $C^{1,\alpha}$ estimates.
+Fix one patch $U_{k}$ and define $v_{k}=\widetilde{u}\circ\Psi_{k}^{-1}$ on $B_{1}^{+}$.
+Then $v_{k}$ solves a uniformly elliptic equation in divergence form with $C^{\alpha}$ coefficients and boundary data $v_{k}=\varphi_{k}$ on $\partial B_{1}^{+}$.
+By the Euclidean boundary $C^{1,\alpha}$ estimate (Gilbarg–Trudinger [5]), we have
+$\|v_{k}\|_{C^{1,\alpha}(\overline{B_{1/2}^{+}})}\leq C\Big(\|v_{k}\|_{L^{\infty}(B_{1}^{+})}+\|\widetilde{f}\|_{C^{\alpha}(B_{1}^{+})}+\|\varphi_{k}\|_{C^{1,\alpha}(\partial B_{1}^{+})}\Big).$
+Step 4: Patch local estimates to the whole domain. Multiplying by $\eta_{k}$ and summing over $k$, we obtain a global estimate on $\Omega$:
+$\|\widetilde{u}\|_{C^{1,\alpha}(\overline{\Omega})}\leq C\Big(\|\widetilde{u}\|_{L^{\infty}(\Omega)}+\|\widetilde{f}\|_{C^{\alpha}(\Omega)}+\|\widetilde{\varphi}\|_{C^{1,\alpha}(\partial\Omega)}\Big).$
+Using the uniform $C^{1,\alpha}$ control of the coordinate chart $\Phi$, we obtain the desired estimate.
+4. Proof of Theorem 1.1
+To prove Theorem 1.1, we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one.
+We first prove
+Theorem 4.1.
+Let $B_{\delta}(x_{j})$ be any ball mentioned above, then there exists a constant $C=C(n,\lambda,i)$ such that
+[…]
+We use Adams’ method to obtain an equivalent version of Theorem 4.1:
 Theorem 4.2.
@@
 […]
-The proof of Theorem 3.3 is exactly the same as that of Theorem 3.1 in [8] which is one of the main results.
+With the help of Green’s function estimates in the last section, the proof of Theorem 4.2 is exactly the same as that of Theorem 3.1 in [10] which is one of the main results.
 The parameters $\lambda,i$ do not explicitly appear in the estimates, but will determine $Q$ and $r_{H}(\alpha,0,Q)$, which gives the bounds of $\delta$.
@@
 This is a constant only depending on the initial geometry of the manifold.
+Here we make some remarks about our alternative proof.
+Remark 4.1.
+To establish global inequalities from local ones, we choose appropriate cutoff functions $\psi_{j}$ such that $\nabla(\psi_{j}^{2}u)||_{L^{n}}\leq 1$.
+Thus the scale parameter $\delta$ actually depends on $\tau$.
+On the other hand, in the Green’s function estimates, we make use of both $W^{1,p}$ bounds of $g$ and $W^{2,p}$ bounds of $x^{i}$.
+There we need to choose proper value of $p$, hence $\alpha$.
+This can be done since the harmonic coordinates exists for arbitrary chosen $\delta$’s and $\alpha$’s.
+In other words, our inequalities can only be proved in specific harmonic coordinates.
+Remark 4.2.
+There is another version of harmonic coordinates one can use to prove our main theorem.
+In [1], it reads as a collection of finitely overlapping charts $\{U_{j},F_{j}\}$, where $F_{j}:U_{j}\to\mathbb{R}^{n}$ is of the form $U_{j}=F_{j}^{-1}(B_{r_{H}}(0))$.
+In such choice of harmonic coordinates, the a priori estimate of $F(x,y)$ will be significantly simplified since the Dirichlet problem is taken on Euclidean balls whose boundary is smooth and can be uniformly finitely covered.
+On the other hand, the domains $\{U_{j}\}$ share the same Gromov’s covering properties.
+Noticing that $\textit{Vol}(U_{j})=\int_{U_{j}}dV=\int_{B_{r_{H}}(0)}\sqrt{|g|}dx\sim C(n,Q)r_{H}^{n}$, it is still possible to perform the local to global argument using this collection of charts, instead of geodesic balls.

### v3

@@
 Abstract.
-The main purpose of this short note, on the one hand, is to clarify some part of the proof of Theorem 1.3 in [10] using the level set argument in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones by using Gromov’s covering lemma.
-Moreover, we show that the Green functions for the Dirichlet problem on any small geodesic balls (or harmonic coordinate charts) on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls (or the charts).
-This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set and thus rigorize the proof of Theorem 1.3 in [10].
+The main purpose of this short note, on the one hand, is to rigorize some part of the proof of Theorem 1.3 in [11] using the level set argument in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones by using Gromov’s covering lemma.
+This result holds on Riemannian manifolds with positive injective radius and Ricci lower bound as stated in [11].
+Moreover, we show that the Green functions for the Dirichlet problem on any small geodesic balls (or harmonic coordinate charts) on Riemannian manifolds with additional condition that the sectional curvature is bounded from above have uniform estimates independent of the locations of the balls (or the charts).
+This will give another proof on such manifolds the critical Trudinger-Moser inequality.
 This latter result of Green’s function estimates is of its independent interest and can be found useful elsewhere.
@@
 $\textit{inj}(M,g)\geq i>0$ for some constant $i>0$.
-Then for every $\tau>0$, there exists a constant $C=C(n,\tau,M)$ such that
+Then there exists a positive constant $\epsilon=\epsilon(\lambda,i,n)=O(\frac{1}{r_{H}})$ in the reverse order of the harmonic radius $r_{H}$ such that for every $\tau>\epsilon$, there exists a constant $C=C(n,\tau,M)$ such that
 […]
 Moreover, $\alpha_{n}$ is sharp.
-In the proof of Theorem 1.1 some clarifications need to be made about the uniformity of the upper bound of the Trudinger-Moser inequality in terms of the bounded domains of the level sets involved.
-In this note, we will then simply give another simple proof by only establishing the uniformity of the bounds of the Trudinger-Moser inequality on geodesic balls within the manifolds (still going through the same proof of Theorem 3.1 in [10] on the geodesic balls or simply applying Theorem 3.1 in [10] on geodesic balls where the Green functions have uniform estimates) together with a Gromov’s covering lemma (similar to the proof of subcritical case in [10]).
-The proof of Theorem 3.1 in [10] remains valid for geodesic balls with uniform constants for the Trudinger-Moser inequalities for such balls independent of the location of the balls.
-The main result of this note is that the Green functions for the Dirichlet problem on any small geodesic balls on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls.
-This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set.
-More precisely, in the proof of Theorem 1.3 in [10], the level set $\Omega(u)$ can be covered by a finite number of small geodesic balls (or harmonic coordinate charts) and the number of such balls (or charts) is uniformly bounded.
-Therefore, the same level argument from local inequalities to the global ones in [10] are justified.
+In the proof of Theorem 1.1 needs to be rigorized about the uniformity of the upper bound of the critical Trudinger-Moser inequality in terms of the bounded domains of the level sets involved.
+In this note, we will then simply give another simple proof by only establishing the uniformity of the bounds of the critical Trudinger-Moser inequality on geodesic balls within the manifolds by modifying the ideas developed in the earlier work by Lam, Lu and Zhang [10].
+In fact, we will achieve this by using a scaling argument and the asymptotic estimates for the supremum of the subcritical Trudinger-Moser functional established in [10].
+As a consequence of this, we can prove the uniform boundedness of the critical Trudinger-Moser inequality on the level set of the functions under consideration in [11].
+This thus rigorizes the proof of Theorem 1.1 (namely, Theorem 1.3 in [11]) by covering the level set by the geodesic balls.
+Therefore, the same level argument from local inequalities to the global ones used in [11] are justified.
+We make use of the equivalence between the subcritical and critical Moser-Trudinger inequalties on geodesic balls.
+This idea was originally in [10] and already used in [11] to derive subcritical inequalities from critical inequalities.
+Meanwhile in this note, we will prove in the opposite direction, i.e.
+from subcritical inequalities to critical inequalities.
+The second main result of this note is to establish that Green functions for the Dirichlet problem on any small geodesic balls on Riemannian manifolds with positive injective radius, Ricci curvature lower bounded and sectional curvature upper bounded have uniform estimates independent of the locations of the balls.
+We believe that this result is of its independent interest.
+More precisely,
+Theorem 1.2.
+Let $(M,g)$ be a complete noncompact Riemannian manifold whose Ricci curvature is lower bounded and sectional curvature is upper bounded, i.e.
+$\textit{Ric}\geq\lambda g$ and $K\leq a^{2}$ for some constant $\lambda,a\in\mathbb{R}$.
+Moreover, we assume that its injectivity radius has a lower bound, i.e.
+$\textit{inj}(M,g)\geq i>0$ for some constant $i>0$.
+For any $Q>1$ and $0<\alpha<1$, denote the geodesic ball as $B_{\delta}(x)$ with $\delta\leq r_{H}(n,i,\lambda,\alpha)$, where $r_{H}(n,i,\lambda,\alpha)$ is the harmonic radius.
+Then the Dirichlet Green’s function over $B_{\delta}(x)$ satisfies the following estimate:
+$|\nabla_{y}G(x,y)|\leq\frac{1}{\omega_{n-1}}d(x,y)^{1-n}(1+Cd(x,y)),$
+where $C=C(n,\lambda,Q,\delta)$.
+As a byproduct, this will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set, within an extra sectional curvature assumption by using the uniform estimates of the Green functions.
 Since the main purpose of this short note is to rigorize the proof of Theorem 1.3 in [11] and in the meantime to give an alternative proof from the level set argument passing from the local inequalities to global one, we have chosen not to give a comprehensive account of the subject (see e.g.
@@
     for any $i\neq j$, $B_{r/2}(x_{i})\cap B_{r/2}(x_{j})=\emptyset$.
-On the other hand, lower bounds for both Ricci curvature and injectivity radius guarantee the existence of $C^{0,\alpha}$ harmonic coordinates.
+[…]
+$\begin{cases}\Delta x^{i}(Q)=0,\ \ \ \textit{for}\ Q\in B_{\rho}(P);\\ x^{i}(P)=0,\frac{\partial x^{i}}{\partial\widetilde{x}^{j}}(P)=\delta^{i}_{j};\\ x^{i}(Q)=\widetilde{x}^{i}(Q)\ \ \ \textit{for}\ Q\in\partial B_{\rho}(P).\end{cases}$
+Its solution $x^{i}$ is called harmonic coordinate for $B_{\rho}(P)$.
+Moreover, lower bounds for both Ricci curvature and injectivity radius guarantee the existence of $C^{0,\alpha}$ harmonic coordinates.
 This is due to Anderson-Cheeger [1] (see also Hebey [6]) and to be precise:
@@
 Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy $\textit{Ric}\geq\lambda g$ and $\textit{inj}(M,g)\geq i>0$.
-Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in\Omega$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
+Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in M$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
 Combining Lemma 2.1 and Lemma 2.2, we obtain a sequence of balls $\{B_{\delta}(x_{j})\}$ satisfies following:
@@
 […]
-Similar embedding also appears for the coordinate functions $x^{i},i=1,\dots,n$ and we have $W^{2,p}$ uniform bounds.
-3. Green’s function estimate
+Similar embedding also appears for the coordinate functions $x^{i},i=1,\dots,n$. We have $W^{2,p}$ , and hence $C^{1,\alpha}$ uniform bounds.
+3. A path from subcritical inequalities to critical ones: Proof of Theorem 1.1
+This section gives a proof of Theorem 1.1 by establishing uniform critical Moser-Trudinger inequalities on geodesic balls from subcritical ones.
+By a covering lemma, we can establish the global ones either by using the level set argument as done in [11] or the covering lemma.
+We first recall some notations.
+Notation.
+Let $(M,g)$ be an $n$-dimensional Riemannian manifold, $n\geq 2$. Fix $\tau>0$ and define
+$\left\lVert u\right\rVert_{1,\tau}:=\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\tau\left\lVert u\right\rVert_{L^{n}(M)}.$
+Let
+$\phi_{n}(t):=e^{t}-\sum_{k=0}^{n-2}\frac{t^{k}}{k!},\qquad\alpha_{n}:=n\Bigl(\frac{n\pi^{n/2}}{\Gamma(\frac{n}{2}+1)}\Bigr)^{\!\!\frac{1}{n-1}}=n\,\omega_{n-1}^{1/(n-1)}.$
+We first quote a result from [10]. The following theorem provides the lower and upper bounds asymptotically for the supremum of the subcritical Trudinger-Moser inequality in $\mathbb{R}^{n}$.
+Theorem 3.1.
+Let $n\geq 2$ and $0\leq\alpha<\alpha_{n}.$ Denote
+$AT\left(\alpha,\beta\right)=\sup_{\left\|\nabla u\right\|_{n}\leq 1}\frac{1}{\left\|u\right\|_{n}^{n-\beta}}\int_{\mathbb{R}^{n}}\phi_{n}\left(\alpha\left(1-\frac{\beta}{n}\right)\left|u\right|^{\frac{n}{n-1}}\right)\frac{dx}{\left|x\right|^{\beta}}.$
+Then there exist positive constants $c=c\left(n,\beta\right)$ and $C=C\left(n,\beta\right)$ such that when $\alpha$ is close enough to $\alpha_{N}:$
+(3.1) $\frac{c\left(n,\beta\right)}{\left(1-\left(\frac{\alpha}{\alpha_{n}}\right)^{n-1}\right)^{\left(n-\beta\right)/}}\leq AT\left(\alpha,\beta\right)\leq\frac{C\left(n,\beta\right)}{\left(1-\left(\frac{\alpha}{\alpha_{n}}\right)^{n-1}\right)^{\left(n-\beta\right)/n}}.$
+Moreover, the constant $\alpha_{N}$ is sharp in the sense that $AT\left(\alpha_{N},\beta\right)=\infty.$
+Theorem 3.2 (Uniform local estimate).
+Given a complete and noncompact Riemannian manifold $M$. Assume there exist $\delta_{0}>0$ and $\Lambda\geq 1$ such that for every $p\in M$ there is a harmonic coordinate chart
+$\varphi_{p}:B_{\delta_{0}}(p)\longrightarrow\mathbb{R}^{n}$
+whose image contains a Euclidean ball $B_{r}(0)$ with $r\in[c_{1}\delta_{0},c_{2}\delta_{0}]$, and such that on $B_{\delta_{0}}(p)$
+$\Lambda^{-1}\left\lvert\xi\right\rvert^{2}\leq g^{ij}(x)\xi_{i}\xi_{j}\leq\Lambda\left\lvert\xi\right\rvert^{2},\qquad\Lambda^{-1}\leq\sqrt{\det(g_{ij}(x))}\leq\Lambda.$
+Then there exists $\delta\in(0,\delta_{0}]$ and $C<\infty$, depending only on $(n,\tau,\Lambda)$, such that for every $u\in W^{1,n}(M)$ with $\left\lVert u\right\rVert_{1,\tau}\leq 1$,
+$\sup_{p\in M}\int_{B_{\delta/2}(p)}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C.$
+Proof.
+Fix $\delta\in(0,\delta_{0}]$ (chosen at the end depending only on $(n,\tau,\Lambda)$) and fix $p\in M$.
+Let $\varphi:=\varphi_{p}$ be the harmonic chart on $B_{\delta}(p)$.
+Write $B_{r}:=B_{r}(0)\subset\varphi(B_{\delta}(p))$ with $r\sim\delta$ (constants depending only on $\Lambda$), and define
+$\widetilde{u}(x):=u(\varphi^{-1}(x)),\qquad x\in B_{r}.$
+By the metric and Jacobian comparability in harmonic coordinates, Euclidean and Riemannian $L^{n}$-norms on $B_{\delta}(p)$ and $B_{r}$ are comparable up to constants depending only on $\Lambda$.
+Step 1: Cutoff (localization). Choose $\eta\in C_{c}^{\infty}(B_{r})$ such that
+$\eta\equiv 1\ \text{on }B_{r/2},\qquad 0\leq\eta\leq 1,\qquad\left\lvert\nabla\eta\right\rvert\leq\frac{C}{r}.$
+For a scaling parameter $a>0$ (to be chosen), set
+$v(x):=a\,\widetilde{u}(x)\quad(x\in B_{r}),\qquad w:=\eta v,$
+and extend $w$ by 0 to $\mathbb{R}^{n}$. Then $w\in W^{1,n}(\mathbb{R}^{n})$ and $w=v$ on $B_{r/2}$, hence
+(3.2) $\int_{B_{r/2}}\phi_{n}\!\left(\alpha_{n}\left\lvert v\right\rvert^{\frac{n}{n-1}}\right)\,dx=\int_{B_{r/2}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx\leq\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx.$
+Step 2: Gradient control for the cutoff. Since $\nabla w=\eta\nabla v+v\nabla\eta$,
+$\left\lVert\nabla w\right\rVert_{L^{n}(\mathbb{R}^{n})}\leq C\left(\left\lVert\nabla v\right\rVert_{L^{n}(B_{r})}+\frac{1}{r}\left\lVert v\right\rVert_{L^{n}(B_{r})}\right).$
+Using $v=a\widetilde{u}$, $r\sim\delta$, and harmonic coordinate comparability, we obtain
+(3.3) $\left\lVert\nabla w\right\rVert_{L^{n}(\mathbb{R}^{n})}\leq C_{1}a\left(\left\lVert\nabla u\right\rVert_{L^{n}(B_{\delta}(p))}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(B_{\delta}(p))}\right)\leq C_{1}a\left(\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}\right).$
+Step 3: Choose $a$ uniformly so that $\left\lVert\nabla w\right\rVert\leq\theta$. Assume $\left\lVert u\right\rVert_{1,\tau}\leq 1$, i.e.
+$\tau\left\lVert u\right\rVert_{L^{n}(M)}+\left\lVert\nabla u\right\rVert_{L^{n}(M)}\leq 1.$
+Fix a margin $\theta:=\frac{1}{2}$ and choose
+(3.4) $a:=\frac{1}{2C_{1}\left(\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}\right)}.$
+Then (3.3) gives
+(3.5) $\left\lVert\nabla w\right\rVert_{L^{n}(\mathbb{R}^{n})}\leq\theta=\frac{1}{2}.$
+(Here $a$ depends on $u$, but the final bound will be uniform because $a^{n}\left\lVert u\right\rVert_{L^{n}}^{n}$ is uniformly controlled by $\left\lVert u\right\rVert_{1,\tau}\leq 1$.)
+Step 4: Apply the asymptotic estimates for the subcritical Trudinger-Moser inequality at a fixed subcritical parameter. Define
+[…]
+Then
+$\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}=\alpha_{n}\theta^{\frac{n}{n-1}}\left\lvert z\right\rvert^{\frac{n}{n-1}}=:\alpha\,\left\lvert z\right\rvert^{\frac{n}{n-1}},\qquad\alpha:=\alpha_{n}\theta^{\frac{n}{n-1}}<\alpha_{n},$
+where $\alpha$ is a fixed subcritical constant depending only on $n$. By the definition of $\mathrm{AT}(\alpha,0)$,
+(3.6) $\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx=\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha\left\lvert z\right\rvert^{\frac{n}{n-1}}\right)\,dx\leq\mathrm{AT}(\alpha,0)\,\left\lVert z\right\rVert_{L^{n}(\mathbb{R}^{n})}^{n}=\mathrm{AT}(\alpha,0)\,\theta^{-n}\,\left\lVert w\right\rVert_{L^{n}(\mathbb{R}^{n})}^{n}.$
+Since $\alpha$ is fixed and $\alpha<\alpha_{n}$, $\mathrm{AT}(\alpha,0)$ is a finite constant depending only on $n$.
+Step 5: Uniform control of $\left\lVert w\right\rVert_{L^{n}}$ and conclusion.
+Because $0\leq\eta\leq 1$ and $w=\eta v$ on $B_{r}$,
+$\left\lVert w\right\rVert_{L^{n}(\mathbb{R}^{n})}\leq\left\lVert v\right\rVert_{L^{n}(B_{r})}=a\,\left\lVert\widetilde{u}\right\rVert_{L^{n}(B_{r})}\leq C_{2}\,a\,\left\lVert u\right\rVert_{L^{n}(B_{\delta}(p))}\leq C_{2}\,a\,\left\lVert u\right\rVert_{L^{n}(M)}.$
+Substituting into (3.6) yields
+$\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx\leq C_{3}(n)\,a^{n}\,\left\lVert u\right\rVert_{L^{n}(M)}^{n}.$
+Now use the definition (3.4).
+Since
+$\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}\geq\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)},$
+we obtain
+$a=\frac{1}{2C_{1}\left(\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}\right)}\leq\frac{\delta}{2C_{1}}\,\frac{1}{\left\lVert u\right\rVert_{L^{n}(M)}}.$
+Therefore
+$a^{n}\,\left\lVert u\right\rVert_{L^{n}(M)}^{n}\leq\left(\frac{\delta}{2C_{1}}\right)^{n},$
+and we conclude
+$\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx\leq C_{4}(n,\Lambda)\,\delta^{n}.$
+Combining with (3.2) gives the Euclidean local bound on $B_{r/2}$. Finally, pulling back by $\varphi$ and using Jacobian comparability yields
+$\int_{B_{\delta/2}(p)}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C(n,\tau,\Lambda),$
+with $C$ independent of $u$ and $p$ (after fixing $\delta$ depending only on $(n,\tau,\Lambda)$). This proves the theorem. ∎
+Next we need to putting together the local estimates over balls to get the estimates over $M$.
+Lemma 3.1 (Cutoff bounds).
+Let $\{x_{j}\}$ be as in Lemma 2.2.
+For each $j$ choose $\psi_{j}\in C_{c}^{\infty}(B_{\delta}(x_{j}))$ such that $0\leq\psi_{j}\leq 1$, $\psi_{j}\equiv 1$ on $B_{\delta/2}(x_{j})$, and $\left\lvert\nabla\psi_{j}\right\rvert\leq 4/\delta$.
+Then $\left\lvert\nabla(\psi_{j}^{2})\right\rvert\leq 8\psi_{j}/\delta$ and for every $u\in W^{1,n}(M)$ and every $\widetilde{\tau}\geq 0$,
+(3.7) $\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}\leq\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\Bigl(\widetilde{\tau}+\frac{8}{\delta}\Bigr)\left\lVert u\right\rVert_{L^{n}(M)}.$
+Moreover,
+(3.8) $\sum_{j}\int_{M}\psi_{j}^{2n}\left\lvert\nabla u\right\rvert^{n}\,dV_{g}\leq N\int_{M}\left\lvert\nabla u\right\rvert^{n}\,dV_{g},\qquad\sum_{j}\int_{M}\left\lvert\nabla(\psi_{j}^{2})\right\rvert^{n}\left\lvert u\right\rvert^{n}\,dV_{g}\leq\frac{8^{n}}{\delta^{n}}\,N\int_{M}\left\lvert u\right\rvert^{n}\,dV_{g}.$
+Proof.
+For (3.7), compute
+$\nabla(\psi_{j}^{2}u)=\psi_{j}^{2}\nabla u+u\,\nabla(\psi_{j}^{2}),$
+hence by the triangle inequality in $L^{n}$,
+$\left\lVert\nabla(\psi_{j}^{2}u)\right\rVert_{L^{n}(M)}\leq\left\lVert\psi_{j}^{2}\nabla u\right\rVert_{L^{n}(M)}+\left\lVert u\,\nabla(\psi_{j}^{2})\right\rVert_{L^{n}(M)}\leq\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{8}{\delta}\left\lVert\psi_{j}u\right\rVert_{L^{n}(M)}\leq\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{8}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}.$
+Also $\left\lVert\psi_{j}^{2}u\right\rVert_{L^{n}}\leq\left\lVert u\right\rVert_{L^{n}}$, so adding $\widetilde{\tau}\left\lVert\psi_{j}^{2}u\right\rVert_{L^{n}}$ gives (3.7). The implication
+$\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}\leq 1$
+follows immediately from $\tau\geq\widetilde{\tau}+\frac{8}{\delta}$ and $\left\lVert u\right\rVert_{1,\tau}\leq 1$.
+Theorem 3.3 (Global Moser–Trudinger from the local estimate).
+2.2 Assume the hypotheses of Lemma  2.2. Fix $\delta\in(0,r_{H}]$ and assume the following local estimate holds:
+(Local MT at scale $\delta$).
+There exist $\widetilde{\tau}\geq 0$ and $C_{\mathrm{loc}}>0$ (depending only on $n$ and the geometric data) such that for every $p\in M$ and every $f\in W^{1,n}(M)$ supported in $B_{\delta}(p)$ with $\left\lVert f\right\rVert_{1,\widetilde{\tau}}\leq 1$,
+(3.9) $\int_{B_{\delta/2}(p)}\phi_{n}\!\left(\alpha_{n}\left\lvert f\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C_{\mathrm{loc}}.$
+Then for every $\tau\geq\widetilde{\tau}+\frac{8}{\delta}$ there exists $C_{\mathrm{glob}}<\infty$ depending only on $(n,\tau)$ and the geometric data such that for all $u\in W^{1,n}(M)$ with $\left\lVert u\right\rVert_{1,\tau}\leq 1$,
+(3.10) $\int_{M}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C_{\mathrm{glob}}.$
+Proof.
+Let $\{x_{j}\}$ and $\{\psi_{j}\}$ be as in Lemmas  2.2 and 3.1. Since $M=\bigcup_{j}B_{\delta/2}(x_{j})$ and $\psi_{j}\equiv 1$ on $B_{\delta/2}(x_{j})$, we have pointwise on $B_{\delta/2}(x_{j})$:
+$\left\lvert u\right\rvert=\left\lvert\psi_{j}^{2}u\right\rvert$.
+Using that $\phi_{n}$ is increasing on $0,\infty)$, we obtain
+(3.11) $\int_{M}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq\sum_{j}\int_{B_{\delta/2}(x_{j})}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}=\sum_{j}\int_{B_{\delta/2}(x_{j})}\phi_{n}\!\left(\alpha_{n}\left\lvert\psi_{j}^{2}u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}.$
+Apply the local estimate with correct scaling.
+Assume $\tau\geq\widetilde{\tau}+\frac{8}{\delta}$ and $\left\lVert u\right\rVert_{1,\tau}\leq 1$.
+Then Lemma 3.1 gives $\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}\leq 1$ for all $j$, and $\psi_{j}^{2}u$ is supported in $B_{\delta}(x_{j})$.
+Thus we may apply (3.9) with $p=x_{j}$ and $f=\psi_{j}^{2}u$ to obtain
+$\int_{B_{\delta/2}(x_{j})}\phi_{n}\!\left(\alpha_{n}\left\lvert\psi_{j}^{2}u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C_{\mathrm{loc}}.$
+To globalize, we use the standard homogeneous consequence of the local estimate:
+Homogeneous local bound.
+There exists $C^{\prime}_{\mathrm{loc}}>0$ such that for all $p\in M$ and all $f$ supported in $B_{\delta}(p)$ and satisfying $\left\lVert f\right\rVert_{1,\widetilde{\tau}}\leq 1$,
+(3.12) $\int_{B_{\delta/2}(p)}\phi_{n}\!\left(\alpha_{n}\left\lvert f\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C^{\prime}_{\mathrm{loc}}\;\ \left\lVert f\right\rVert_{1,\widetilde{\tau}}^{\,n}.$
+Applying (3.12) to $f=\psi_{j}^{2}u$ and summing in (3.11) gives
+$\int_{M}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C^{\prime}_{\mathrm{loc}}\sum_{j}\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}^{\,n}.$
+Estimate the sum of localized norms using bounded overlap. Using
+$(A+B)^{n}\leq 2^{n-1}(A^{n}+B^{n}),$
+the product rule, and $\psi_{j}\leq 1$,
+$\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}^{\,n}\leq C(n)\int_{M}\psi_{j}^{2n}\left\lvert\nabla u\right\rvert^{n}+\left\lvert\nabla(\psi_{j}^{2})\right\rvert^{n}\left\lvert u\right\rvert^{n}+\widetilde{\tau}^{n}\psi_{j}^{2n}\left\lvert u\right\rvert^{n}\,dV_{g}.$
+Summing over $j$ and using Lemma 3.1 and $\sum_{j}\psi_{j}^{2n}\leq N$) yields
+$\sum_{j}\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}^{\,n}\leq C(n)N\int_{M}\left\lvert\nabla u\right\rvert^{n}\,dV_{g}+C(n)N\left(\frac{8^{n}}{\delta^{n}}+\widetilde{\tau}^{n}\right)\int_{M}\left\lvert u\right\rvert^{n}\,dV_{g}.$
+If $\left\lVert u\right\rVert_{1,\tau}\leq 1$ then $\int_{M}\left\lvert\nabla u\right\rvert^{n}\leq 1$ and $\int_{M}\left\lvert u\right\rvert^{n}\leq\tau^{-n}$. Therefore
+$\int_{M}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C^{\prime}_{\mathrm{loc}}\,C(n)N\left(1+\left(\frac{8^{n}}{\delta^{n}}+\widetilde{\tau}^{n}\right)\tau^{-n}\right)=:C_{\mathrm{glob}},$
+which is finite and depends only on $(n,\tau)$ and the geometric data. This proves (3.10). ∎
+4. Green’s function estimate: Proof of Theorem 1.2
 For $u\in W^{1,n}_{0}(B_{\delta}(x_{j}))$, we recall the pointwise representation formula of $u(x)$ in terms of the Dirichlet Green’s function:
@@
 $G(x,y)=H(x,y)+\sum_{l=1}^{k}\int_{B_{\delta}(x_{j})}\Gamma_{l}(x,z)H(z,y)dV_{g}(z)+F(x,y),$
-where $H(x,y)=\frac{1}{(n-2)\omega_{n-1}}d(x,y)^{2-n}f(d(x,y))$ is the fundamental solution of Laplacian multiplied by a smooth cutoff function $f$, which is identically equal to 1 near diagonal and vanishes when $d(x,y)\geq\textit{inj}(M,g)$.
+where $H(x,y)=\frac{1}{(n-2)\omega_{n-1}}d(x,y)^{2-n}f(d(x,y))$ is the fundamental solution of Laplacian multiplied by a smooth cutoff function $f$, which is identically equal to 1 when $d(x,y)<\frac{\delta}{2}$ and vanishes when $d(x,y)\geq\delta$.
+Moreover, $f$ can be chosen such that $|f^{\prime}|\leq\frac{C}{\delta},|f^{\prime\prime}|\leq\frac{C}{\delta^{2}}$ for some uniform constant $C>0$.
 $\Gamma(x,y)=\Gamma_{1}(x,y)=-\Delta_{y}H(x,y)$ and we inductively define
@@
 (4.2) $|\nabla_{y}G(x,y)|\leq\frac{1}{\omega_{n-1}}d(x,y)^{1-n}(1+Cd(x,y)),$
-where $C=C(n,\lambda,Q,\lambda,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
+where $C=C(n,\lambda,Q,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
 $\int_{B_{\delta}(x_{j})}\Gamma(x,z)H(z,y)dV_{g}(z).$
-From the asymptotic behavior near singularity, we need to prove the following lemma:
+For the asymptotic behavior near singularity, we need to prove the following lemma:
 Lemma 4.1.
@@
 (4.3) $\int_{B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
-(3.4) $=$ $\left(\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}+\int_{(B_{3\rho}(y)\setminus B_{\rho}(x))\cap B_{\delta}(x_{j})}+\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}\right)d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
-It is well known that in polar coordinates
-$dV=r^{n-1}\left(1-\frac{\textit{Ric}(\theta)}{3}r^{2}+o(r^{2})\right)d\theta dr.$
-When $\delta$ is small, providing Ricci curvature lower bounded, the first term of the right hand side,
+(4.4) $=$ $\left(\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}+\int_{(B_{3\rho}(y)\setminus B_{\rho}(x))\cap B_{\delta}(x_{j})}+\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}\right)d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z).$
+The first term of the right hand side can be estimated as
 $\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
 $\leq$ $\rho^{\beta-n}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}dV_{g}(z),$
-$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
-$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\alpha+\beta-n},$
-Similarly we can estimate the second term of the right hand side of (3.3). For the third term, we have
+$\leq$ $C(\lambda,\delta,Q)\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
+$\leq$ $C(\lambda,\delta,Q)\omega_{n-1}\rho^{\alpha+\beta-n},$
+where from the second line to the third line we use the property that in harmonic coordinate, $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ as bilinear form.
+Hence integral can be transferred to normal coordinate and further to polar coordinate.
+Meanwhile it is worth to emphasize that we omit the transition function between harmonic coordinate charts centered at $x_{j}$ and centered at $x$, since it is a $C^{1,\alpha}$ function and can just be treated as a uniform constant.
+We will also apply such treatment in future estimates.
+Similarly we can estimate the second term of the right hand side of (4.3).
+For the third term, we have
 (4.5) $\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
 (4.6) $\leq$ $C_{n}\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(z,y)^{\alpha+\beta-2n}dV_{g}(z),$
-(3.7) $\leq$ $C_{n}C(\lambda,\delta)\omega_{n-1}\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$
+(4.7) $\leq$ $C(n,\lambda,\delta,q)\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$
 (4.8) $\leq$ $C_{n,\alpha,\beta,\lambda,\delta}\omega_{n-1}((3\rho)^{\alpha+\beta-n}-\delta^{\alpha+\beta-n}),$
@@
 It is easy to verify that in polar coordinates,
-[…]
-where $r=d(x,y)$ and $|\widetilde{g}|=\textit{det}(\widetilde{g}_{ij})$ denotes the determinant of the corresponding metric tensor.
-By carefully choosing the smooth cut-off function $f$, both the first two terms $|(n-3)f^{\prime}-rf^{\prime\prime}|$ and the coefficient of the third term $((n-2)f-rf^{\prime})$ can be uniformly bounded by $C(n,\delta,i)$.
-With the help of Lemma 3.1, we only need to consider the integral of the form
-$\int_{B_{\delta}(x_{j})}|\nabla_{z}\ln\sqrt{|\widetilde{g}|}|(z)H(z,y)dV_{z}.$
-An easy change of variablse from polar coordinates to harmonic coordinates shows
-$\widetilde{g}_{ij}=g_{kl}\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\frac{\partial x^{l}}{\partial\widetilde{x}^{j}},$
-where $\{x^{k}\}$ and $\{\widetilde{x}^{i}\}$ denotes the coordinate variables respectively.
-For simplicity we denote the transition matrix $A=\left(\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\right)$ then
-$|\nabla\ln\sqrt{|\widetilde{g}|}|\leq\frac{|\nabla|g||}{2|g|}+\frac{|\nabla|A||}{|A|}.$
-Since in harmonic coordinates $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$, we have
-$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|g||}{|g|}(z)H(z,y)dV_{z}\right|$ $\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$
-Similar to the argument in the proof of Lemma 3.1, the integral can be splitted into 3 pieces.
-Here we just check the part over $B_{\rho}(x)\cap B_{\delta}(x_{j})$, the other two parts can be treated in a similar manner:
-$\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$ $\leq\frac{C(n)}{d(x,y)^{n-2}}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}|\nabla g|dV_{z};$
-[…]
-$\leq\frac{C(n,Q)}{d(x,y)^{n-2-\frac{n}{p^{\prime}}}}.$
-On the other hand, $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ implies there exists $C(n,Q)>0$ such that $|A|\geq C(n,Q)>0$. Hence we have
-$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|A||}{|A|}(z)H(z,y)dV_{z}\right|\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla|A|(z)|}{d(z,y)^{n-2}}dV_{z}$
-We can estimate the above integral in a similar manner, but this time we make use of the $W^{2,p}$ estimate of the harmonic coordinate function $x^{i},i=1,\dots,n$.
-Indeed, in the derivation of the existence of harmonic coordinates, the $L^{p}$ theory of elliptic operators gives the uniform bound of $x^{i}||_{W^{2,p}}$, where $p$ satisfies $\alpha=1-\frac{n}{p}$.
-Hence there exists $C(n,\lambda,i,Q)$ such that $||\nabla A|_{L^{p}}\leq C(n,\lambda,i,Q)$.
-Therefore from Lemma 3.1, $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a constant $C=C(n,\lambda,i,Q)$.
+$\Delta_{z}H(x,z)=\frac{1}{(n-2)\omega_{n-1}r^{n-1}}\left((n-3)f^{\prime}-rf^{\prime\prime}+((n-2)f-rf^{\prime})\left[\Delta_{z}d(x,z)-\frac{n-1}{d(x,z)}\right]\right),$
+where $r=d(x,z)$.
+By carefully choosing the smooth cut-off function $f$, the first two terms $|(n-3)f^{\prime}-rf^{\prime\prime}|$ can be uniformly bounded by $C(n,\delta,i)$.
+With the help of Lemma 4.1, we need to consider the integral
+$\left|\int_{B_{\delta}(x_{j})}\frac{1}{d(x,z)^{n-1}}((n-2)f-rf^{\prime})\left[\Delta_{z}d(x,z)-\frac{n-1}{d(x,z)}\right]\frac{1}{d(z,y)^{n-2}}dV_{z}\right|.$
+First notice that $|(n-2)f-rf^{\prime}|\leq C(n,\delta,i)$. Applying Bishop comparison theorem and Günther comparison theorem,
+$(n-1)a\cot ad(x,z)\leq\Delta_{d}(x,z)\leq(n-1)\sqrt{-\lambda}\coth\sqrt{-\lambda}d(x,z).$
+Using Taylor expansions for $\cot t$ and $\coth t$, there exists a uniform constant $C=C(n,\delta,i,a,\lambda)$ such that $|\Delta_{z}d(x,z)-\frac{n-1}{d(x,z)}|\leq Cd(x,z)$.
+Therefore from Lemma 4.1, $\Gamma_{2}(x,y)$ is bounded by $d(x,y)^{n-3}$, up to some uniform constant.
+Then we argue inductively and hence conclude that $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a uniform constant (for its derivatives, we leave the proof to readers).
 To obtain the uniform estimate for Green’s function, it suffices to prove the uniform $C^{1}$ estimate of $F(x,y)$.
 Lemma 4.2.
-Let $F(x,y)$ be a $W^{1,2}_{0}(B_{\delta}(x_{j}))$ solution of the Dirichlet problem (3.1), there exists a positive constant $C=C(n,\lambda,i,Q)$ such that $\nabla_{y}F(x,y)||_{\infty}\leq C$.
+Let $F(x,y)$ be a $W^{1,2}_{0}(B_{\delta}(x_{j}))$ solution of the Dirichlet problem (4.1), there exists a positive constant $C=C(n,\lambda,a,i,Q)$ such that $\nabla_{y}F(x,y)||_{\infty}\leq C$.
 Proof.
@@
 Using the uniform $C^{1,\alpha}$ control of the coordinate chart $\Phi$, we obtain the desired estimate.
-4. Proof of Theorem 1.1
-To prove Theorem 1.1, we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one.
+5. Alternative proof of Trudinger-Moser inequalities using the uniform estimates of Green functions
+In this section, we give some remarks about how to prove Theorem 1.1 under additional condition of sectional curvature being upper bounded using the uniform Green function estimates.
+we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one.
 We first prove
@@
 […]
-With the help of Green’s function estimates in the last section, the proof of Theorem 4.2 is exactly the same as that of Theorem 3.1 in [10] which is one of the main results.
+With the help of Green’s function estimates in the previous section, the proof of Theorem 5.2 is exactly the same as that of Theorem 3.1 in [11].
 The parameters $\lambda,i$ do not explicitly appear in the estimates, but will determine $Q$ and $r_{H}(\alpha,0,Q)$, which gives the bounds of $\delta$.
 This shows the constants in the Moser-Trudinger inequality on any geodesic balls only depend on the radius $\delta$ which in turn depends on $\lambda$ and $i$.
-Proof of Theorem 1.1.
-For each $j$, we take cutoff function $\psi_{j}\in C_{0}^{\infty}(B_{\delta}(x_{j}))$ satisfying $0\leq\psi_{j}\leq 1$, $\psi_{j}=1$ on $B_{\frac{\delta}{2}}(x_{j})$ and $|\nabla\psi_{j}|\leq\frac{4}{\delta}$.
-Then $|\nabla\psi_{j}^{2}|=2\psi_{j}|\nabla\psi_{j}|\leq\frac{8}{\delta}\psi_{j}$ and $1\leq\sum_{j}\psi_{j}(x)\leq N$.
-Thus we have
-(4.1) $\int_{M}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g}$
-(4.2) $\leq$ $\sum_{j}\int_{B_{\delta/2}(x_{j})}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g},$
-(4.3) $\leq$ $\sum_{j}\int_{B_{\delta}(x_{j})}\phi_{n}(\alpha_{n}|\psi_{j}^{2}u|^{\frac{n}{n-1}})dV_{g},$
-[…]
-[…]
-(4.6) $\leq$ $C(n,\lambda,i)2^{n}\sum_{j}\int_{M}\psi_{j}^{2n}|\nabla u|^{n}+|\nabla\psi_{j}^{2}|^{n}|u|^{n}dV_{g},$
-(4.7) $\leq$ $C(n,\lambda,i)2^{n}\left(N\int_{M}|\nabla u|^{n}dV_{g}+\frac{8^{n}}{\delta^{n}}N\int_{M}|u|^{n}dV_{g}\right),$
-(4.8) $\leq$ $C(n,\lambda,i)(N+N\frac{8^{n}}{\delta^{n}\tau^{n}}).$
-This is a constant only depending on the initial geometry of the manifold.
-Here we make some remarks about our alternative proof.
-Remark 4.1.
-To establish global inequalities from local ones, we choose appropriate cutoff functions $\psi_{j}$ such that $\nabla(\psi_{j}^{2}u)||_{L^{n}}\leq 1$.
-Thus the scale parameter $\delta$ actually depends on $\tau$.
-On the other hand, in the Green’s function estimates, we make use of both $W^{1,p}$ bounds of $g$ and $W^{2,p}$ bounds of $x^{i}$.
-There we need to choose proper value of $p$, hence $\alpha$.
-This can be done since the harmonic coordinates exists for arbitrary chosen $\delta$’s and $\alpha$’s.
-In other words, our inequalities can only be proved in specific harmonic coordinates.
+Proof of Theorem 1.1 under sectional curvature upper bounded condition then follows from the uniform Green function estimates.
 Remark 5.1.